31
May
2012
13:39

Apologies for the outage affecting all services over the last hour. The root cause appears to be a major instability problem at the LINX, affecting virtually all the UK internet. Our transit providers have temporarily disconnected from there until it is resolved, so service has now been restored.

Posted by Kevin Hones

30
May
2012
13:38

Blusquare 2 and 3 will have no ups backup until 22:00 tonight so racks are running on mains power only. Our core routers, mail servers etc are supported by our own ups within our racks so ADSL services were unaffected by the power outage last night. Sadly the problem is beyond our control and Pulsant assure us that comprehensive testing will be done before the ups systems are reconnected at 22:00 this evening. We do apologise to those CoLo customers who experienced outages last night.

Posted by Phil McGillivray

30
May
2012
13:37

There has been a power failure in Bluesquare 2 and 3 as a result of UPS problems. Pulsant engineers are repairing and at this time power is restored however the UPS is out of service.

Posted by Yolanda

29
May
2012
22:56

Some of our customers will be seeing some issues as a result of core routing problems being experienced by our transit and backhaul provider. Their engineers are investigating the problem. Sorry for any inconvenience you may be experiencing.

Posted by Phil McGillivray

10
May
2012
13:24

All connections have now restored however we are still awaiting an update from BT. Again we do apologise to those customers affected.

Posted by Phil McGillivray

10
May
2012
12:55

BT have raised a fault and are investigating their end. Update due by 13:15. Most lines have reconnected to the backup LNS.

Posted by Phil McGillivray

10
May
2012
12:10

There is a BT fibre link fault – lines are moving to a backup LNS automatically. We apologise to those customers affected by this issue.

Posted by Phil McGillivray

09
May
2012
10:19

Most routes now appear to be working normally, including those to The BBC. If anyone is still experiencing problems, please call support and we will help diagnose.

Posted by Kevin Hones

09
May
2012
10:04

We believe the problem to be at one of our upstream providers connections at the LINX. We will post details when we have them. Most routes have returned, but not all. Please accept our apologies for any inconvenience.

Posted by Kevin Hones

09
May
2012
09:23

We are aware of routing issues upstream of our network. This is affecting connectivity to some large sites such as Google and The BBC. We are investigating the issues further. We apologise for any inconvenience or disruption caused.

Posted by Yolanda

Current Status

    GoodNo current issues

Status Archive